OpenStudy (crashonce):

If A and B are roots of the equation 3x^2 + kx +9=0, then logA^2B - logA=?

OpenStudy (crashonce):

@mathmate

OpenStudy (perl):

does your question say logA^(2B) - logA or log(A^2) * B - log A

OpenStudy (perl):

Applying Vieta's formula to a second degree polynomial (quadratic) P(x)=ax^2 + bx + c the roots r1,r2 of the equation P(x)=0 satisfy r1 + r2 = -b/a r1*r2= c/a

OpenStudy (solomonzelman):

wait, it can also be \(\large\color{black}{\log({\rm A^2B}) - \log({\rm A}) }\)

OpenStudy (crashonce):

^

OpenStudy (perl):

Applying Vieta's formula on your quadratic equation 3x^2 + kx +9 =0 we have A + B = -k/3 A*B = 9/3 = 3 AB = 3 Now we can simplify your log expression: log ( A^2B) - log (A) = log(A^2) + log (B) - log A = 2*log(A) + log (B) - log (A) = log(A) + log(B) but we know AB = 3 log(AB) =log(3) and since log(AB) = log(A) + log(B) ... can you solve it now

OpenStudy (crashonce):

wats the answer, I cant solve it

OpenStudy (crashonce):

k=-9?

OpenStudy (crashonce):

@perl

OpenStudy (perl):

its not asking to solve for k

OpenStudy (crashonce):

@perl the question says to stop at log 3 thaks

OpenStudy (mathmate):

@CrashOnce You have to realize that the question is lacking parentheses, and is ambiguous. It can be one of three interpretations. Perl has suggested two interpretations, and here's the third by @SolomonZelman . You will not evaluate the expression until the question is clear. You can work along the following lines if the third interpretation is valid: or, if \(\large\color{black}{\log({\rm A^2B}) - \log({\rm A}) }\) as @SolomonZelman suggested, we can simplify as \(\large=\color{black}{\log({\rm \frac{A^2B}{A}}) }\) \(\large=\color{black}{\log({\rm AB}) }\) \(\large=...\) Use Vieta's formula to find the value of AB (using hints from @perl )